OpenStudy (anonymous):

Which expression is equivalent to (sin x + 1)(sin x − 1)?

OpenStudy (anonymous):

cos^2x -cos^2 cos^2x+1 cos^2x-1 -cos^2x+1

OpenStudy (mathstudent55):

Can you multiply out the given binomials?

OpenStudy (mathstudent55):

Use FOIL, or the pattern (a + b)(a - b) = a^2 - b^2

OpenStudy (anonymous):

Can you show me how to do that?

OpenStudy (mathstudent55):

Let's use FOIL: \((\sin x + 1)(\sin x - 1) \) \( = \sin^2 x - \sin x + \sin x - 1^2\) \( = \sin^2 x - 1\) Ok?

OpenStudy (anonymous):

Is the answer -cos^2x

OpenStudy (mathstudent55):

Now use the identity: \(\sin^2 x + \cos^2 x = 1\) Subtract \(\sin^2 x \) from both sides. What do you get?

OpenStudy (mathstudent55):

No, but close.

OpenStudy (anonymous):

-cos^2+1?

OpenStudy (mathstudent55):

Instead of guessing, do this: Take this equation, \(\sin^2 x + \cos^2 x = 1\), and subtract \(\sin^2 x\) from both sides. What do you get?

OpenStudy (anonymous):

cos^2x-1?

OpenStudy (mathstudent55):

 Ok?